OpenStudy (anonymous):

Find an expression for the length of an arc subtended by each of the following angles in a circle of radius r. a. 60° b. 270° c. –90°

OpenStudy (anonymous):

I think we need a figure

OpenStudy (anonymous):

there isnt one :(

OpenStudy (anonymous):

OK. So I assume you are talking about angles measured at the origin of the circle, rather than formed on the edge of the circle.

OpenStudy (anonymous):

\[\theta = x\]

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

now that we know how to calculate the circumference of a circle, we can also calculate the length of an arc (which is simply a portion of the circumference). An angle α defined by two radii subtends an arc. Let's take a look at several examples, from which we can identify a pattern. The arc K in each case is shown as a bold curve. The circumference of the circle is C.

OpenStudy (anonymous):

k=1/4 C

OpenStudy (anonymous):

SO T thats a quarter of 360

OpenStudy (anonymous):

OK so the circumference C = 2 pi r for an angle x the fraction f of 360 degrees is x/360 the formula for the arc length L L = (x/360) 2 pi r

OpenStudy (anonymous):

For the example you are talking about, arc length (k) k = 1/4 C = (1/4) 2 pi r. what is r?

OpenStudy (anonymous):

so plug in the 60 in plac of the x

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

radius

OpenStudy (anonymous):

so it would l=60/360*2

OpenStudy (anonymous):

it would be 60/360 times 2 pi r

OpenStudy (anonymous):

right sorry th

OpenStudy (anonymous):

right sorry th

OpenStudy (anonymous):

right sorry th

OpenStudy (anonymous):

right sorry th

OpenStudy (anonymous):

L

OpenStudy (anonymous):

for an angle of 60. for an angle of 90, it would be 90/360 times 2 pi r

OpenStudy (anonymous):

my answer is a fraction right

OpenStudy (anonymous):

yes, if they didn't give you a value for r

OpenStudy (anonymous):

got it U are awESOME

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

@telliott99 ONe first ?? it was 2188 did u round it off to 22